OpenStudy (anonymous):

3+2cos^2(x)=5cos^2(x)+3sin^2(x)

OpenStudy (mertsj):

Replace sin^2x with 1-cos^2x. Get it equal to 0 and factor.

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

3+ 2cos^2x= 5cos^2x+ 3sin^2x lhs= 3+ 2cos^2x = 3( sin^2x+cos^2x) + 2cos^2x =3 sin^2x+3cos^2x + 2cos^2x =5cos^2x+ 3 sin^2x =rhs put sin^2x +cos^2x = 1